OpenStudy (anonymous):

A car passes a landmark on a highway traveling at a constant rate of 55 kilometers per hour. One hour later, a second car passes the same landmark traveling in the same direction at 63 kilometers per hour. How much time after the second car passes the landmark will it overtake the first car? Do not do any rounding.

OpenStudy (anonymous):

@sangya21 @kah.x @mathmath333 @mathstudent55 @uri PLEASE HELP

OpenStudy (anonymous):

not sure but something like this Let t = time for the 2nd car to catch the 1st From the information given, we know that the first car is 55 km ahead of the 2nd car when it passes the marker. write a distance equation, dist = speed * time 63t = 55t + 55 63t - 55t = 55 12t = 55 t = 55/12 t = 4.6 hrs or 4 hrs 36 min

OpenStudy (anonymous):

What about this one Two trains leave the station at the same time, one heading east and the other west. The eastbound train travels 16 miles per hour slower than the westbound train. If the two trains are 680 miles apart after 4 hours, what is the rate of the eastbound train?

OpenStudy (anonymous):

D=RT 680=(X+X-16)*4 680=(2X-16)4 680=8X-64 8X=680+64 8X=744 X=93 X-16= 93-16 EASTBOUND TRAIN =77mph

OpenStudy (anonymous):

Keith drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 9 hours. When Keith drove home, there was no traffic and the trip only took 4 hours. If his average rate was 40 miles per hour faster on the trip home, how far away does Keith live from the mountains?

OpenStudy (anonymous):

do u kno this one @sangya21

OpenStudy (anonymous):

9y = 4(y+40) 9y = 4y +160 5y = 160 y = 32 mph y(return) = 32+40 y(return) = 72 mph 9*32 = 4*72 288=288

OpenStudy (anonymous):

Two trains leave the station at the same time, one heading west and the other east. The westbound train travels at 105 miles per hour. The eastbound train travels at 95 miles per hour. How long will it take for the two trains to be 560 miles apart? Do not do any rounding.

OpenStudy (anonymous):

combined speed = 105+95 = 200mph t = d/speed t = 560/200 t = 2.8hrs = 2hrs 48 mins

OpenStudy (anonymous):

this is my last one Two cars leave towns 800 kilometers apart at the same time and travel toward each other. One car's rate is 16 kilometers per hour less than the other's. If they meet in 4 hours, what is the rate of the slower car?

OpenStudy (anonymous):

@sangya21 do u kno

OpenStudy (anonymous):

D=RT 800=[(R-16)+R]4 800=(2R-16)4 800=8R-64 8R=800+64 8R=864 R=864/8 R= 108 KMH IS THE SPEED OF THE FASTER CAR. 108-16=92KMH IS THE SPEED OF TH SLOWER CAR. PROOF: 800=(108+92)4 800=200*4 800=880

OpenStudy (anonymous):

thankkkk youuu soo muchh @sangya21

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

i have two on inequalities... can u help? @sangya21

OpenStudy (anonymous):

I will try.

OpenStudy (anonymous):

At the state fair, admission at the gate is $11 . In addition, the cost of each ride is $3 . Suppose that Kala will go on x rides. Kala wants the total number of dollars she spends on admission and rides to be at most t . Using the values and variables given, write an inequality describing this.

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

11+3x<=t (less than equal to )

OpenStudy (anonymous):

Kira takes classes at both Westside Community College and Pinewood Community College. At Westside, class fees are $98 per credit hour, and at Pinewood, class fees are $115 per credit hour. Kira is taking a combined total of 19 credit hours at the two schools. Suppose that she is taking w credit hours at Westside. Write an expression for the combined total dollar amount she paid for her class fees.

OpenStudy (anonymous):

thank u ,:)

OpenStudy (anonymous):

98w + 115(19-w) we did this question

OpenStudy (anonymous):

thank u , Alan drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 7 hours. When Alan drove home, there was no traffic and the trip only took 4 hours. If his average rate was 27 miles per hour faster on the trip home, how far away does Alan live from the mountains?

OpenStudy (anonymous):

7y = 4(y+27) solve for y.

OpenStudy (anonymous):

Two trains leave stations 238 miles apart at the same time and travel toward each other. One train travels at 75 miles per hour while the other travels at 95 miles per hour. How long will it take for the two trains to meet? Do not do any rounding.

OpenStudy (anonymous):

238 = (75+95)* t find t

OpenStudy (anonymous):

so 238 = (170)* t then what?

OpenStudy (anonymous):

@sangya21 whats t

OpenStudy (anonymous):

t = time

OpenStudy (anonymous):

what i do next @sangya21

OpenStudy (anonymous):

238/170 = t t = 1.4 t = 1hr and (0.4*60) mins t = 1 hr and 24 mins